# Improving the Language Understanding Capabilities of Large Language Models Using Reinforcement Learning

### v2

@@ 4.3 Proximal Policy Optimization for LLM Fine-tuning on NLU Tasks
We utilize PPO to fine-tune the LLM on NLU tasks, following the training protocol outlined in section 3. The reward function is specifically designed for each NLU task. In this work, we use a simple reward function, where a reward is assigned at the end of the generation based on alignment with the ground truth labels. We use regular expression matching to extract answers from the LLM’s outputs by first locating the text within the ‘<Judgement></Judgement>’ tags. Depending on the task, we then search for task-specific keywords (such as “yes”, “no”, “acceptable”, or “not acceptable”) to identify the answer. These extracted answers are compared with the ground truth to determine the appropriate rewards.
@@ 4.4 Low-Rank Adaptation
To mitigate the computational cost of full-model fine-tuning, we employ LoRA Hu et al. (2021b) during both the supervised fine-tuning and PPO stages. Instead of updating the entire model, we restrict the updates to LoRA layers, which significantly reduces the number of trainable parameters by decomposing the weight matrices into low-rank matrices.
@@ 5.1 Experimental Setup
We trained and evaluated our models on the GLUE Wang et al. (2019), and SuperGLUE Wang et al. (2020) benchmarks. All experiments were conducted using instruction-tuned LLAMA2-7B models Touvron et al. (2023a) [^1] https://huggingface.co/daryl149/llama-2-7b-chat-hf. We perform both single task and multi-task fine-tuning: 1) Single-task Fine-tuning: For each subtask within GLUE, and SuperGLUE, a separate task-specific LoRA module was trained independently. 2) Multi-task Fine-tuning: In the multi-task setting, datasets from different subtasks within each benchmark were combined, and a single LoRA module was trained to handle all tasks simultaneously.
Hyperparameter Settings
For PPO-based fine-tuning, gird search is performed to select the batch size in 4, 8, 12, and 16 for each task. A batch size of 24 was used across all tasks during supervised fine-tuning (SFT). The PPO epochs is set to 4, that is each sampled batch is used for updating the model four times. The initial learning rate for all tasks was set to $9\times 10^{-6}$. We utilized the Adafactor optimizer for PPO training and AdamW for SFT. A cosine annealing learning rate scheduler with a warmup phase was employed, where the learning rate was gradually increased during the first 10% of training steps and then reduced to one-tenth of the initial value by the end of training. We use a rank $r=16$ for the LoRA layers. We trained both PPO and SFT models until convergence on the validation set. The best hyperparameters were selected based on performance on the validation set. The final reported results for the GLUE and SuperGLUE are from their corresponding evaluation server. For evaluation, multinomial sampling with a temperature of 1 was used to generate a single response per data sample. The model generated responses with lengths between 12 and 32 tokens, with the generation process concluding using a special identifier ‘‘</Judgement>’’.
@@ 5.2 Baselines
  • Encoder-only models: We compare our results with encoder-only transformer models, specifically BERT-base (110M parameters) and BERT-large (340M parameters) Devlin et al. (2019).
@@
  • Few-shot prompting: In this setting, the model is provided with both the task-specific prompt and five labeled examples from the training dataset as demonstrations. These examples are provided as reference to guide the model in generating more accurate responses for the input query. Similarly, no task-specific fine-tuning is performed.
@@ 5.3 Results on GLUE Benchmark
In this section, we present our experiments on the GLUE benchmark, comparing the results with encoder-only models such as BERT Devlin et al. (2019). We use the LLAMA2-7B model as the LLM for our evaluations. The baselines include zero-shot prompting and few-shot prompting (5-shot). For fine-tuning methods, we compare both supervised fine-tuning and PPO across single-task and multi-task settings. The results are summarized in Table 1. From the results, we make the following observations.
@@
Table 1: GLUE test results are scored by the evaluation server (GLUE benchmark). Average column indicates the averaged performance across all the datasets excluding the WNLI and AX datasets. F1 scores are reported for QQP and MRPC, Spearman correlations for STS-B, Matthew’s correlations for CoLA, and accuracy scores for the other tasks. Zero-shot prompting refers to prompting with task-specific prompts and an input query, while Few-shot prompting refers to prompting with task-specific prompts, 5 demonstrations, and an input query. PPO stands for proximal policy optimization, and SFT refers to Supervised Fine-tuning. “ST” represents Single-task, while “MT” represents Multi-task. The bolded results indicate the best results, and the underlined results indicate the second-best results.
@@
First, we observed that zero-shot prompting of the LLAMA2-7B model with task-specific prompts consistently underperformed compared to the smaller BERT-base model. LLAMA2-7B struggled notably on simpler tasks like SST-2, which only required classifying sentiment as positive or negative. This underscores the model’s weak language understanding capabilities, with zero-shot prompting proving inadequate compared to BERT-base. Second, few-shot prompting showed improvements over the zero-shot baseline, achieving an average score of 58.7 compared to 46.1, but it still lagged significantly behind the BERT-base model’s score of 79.6. Third, supervised fine-tuning (SFT) using LoRA modules for each task further boosted performance, bringing it closer to BERT’s level with an average score of 78.5, though still slightly behind BERT-base’s 79.6. Fourth, fine-tuning with PPO delivered the best results, achieving an average score of 84.6, surpassing even BERT-large’s 82.1. Moreover, zero-shot and few-shot prompting of LLAMA2-7B displayed a noticeable output imbalance, with a tendency to favor certain classes or values. In contrast, models fine-tuned with PPO showed no significant bias. Fifth, the computational time for PPO is approximately 1.32 times that of SFT, indicating only a marginal increase in computational costs.
@@
Additionally, we compared the results with multi-task training, where a single LoRA module was trained across all datasets using both SFT and PPO to reduce time complexity. We found that SFT on individual tasks outperformed its multi-task fine-tuning counterpart. However, while PPO on multi-task training did not perform as well as PPO on single-task training, it still outperformed BERT-large in average performance, achieving a score of 82.9 compared to BERT-large’s 82.1. These results demonstrate that while single-task fine-tuning yields the best performance, multi-task training with PPO can still achieve competitive results, even surpassing state-of-the-art models like BERT-large. The training curves for PPO is presented in appendix A.
@@
Table 2: SuperGLUE test results are scored by the evaluation server (SuperGLUE benchmark). The experimental data for BERT-large and BERT-large++ are taken from the original SuperGLUE paper Wang et al. (2020). The metrics used in the experiments are as follows: CB: F1 / Acc; MultiRC: F1 / Exact Match; ReCoRD: F1 / Exact Match; AXb: MCC; AXg: Gender parity score / Acc. For the remaining tasks not mentioned, accuracy (Acc) is reported. Average column corresponds to the averaged performance across all the datasets. For tasks with multiple evaluation metrics, we first compute the average of those metrics to obtain a single task score, which is then used in the overall average calculation. The bolded results indicate the best results, and the underlined results indicate the second-best results.
5.4 Results on SuperGLUE Benchmark
@@
We fine-tuned the LLAMA2-7B model using PPO on the SuperGLUE dataset and compared its performance against several baselines, including BERT-large, BERT-large++, and zero-shot and few-shot prompting of LLAMA2-7B. The term “BERT++” refers to a BERT model fine-tuned using the supplementary training on intermediate labeled-data tasks (STILTs) approach Phang et al. (2018), where the model is first fine-tuned on related transfer tasks before being fine-tuned on SuperGLUE tasks. For example, MNLI from the GLUE benchmark Wang et al. (2019) is used as an intermediate task for CB, RTE, and BoolQ Wang et al. (2020). In contrast, our experiments with LLM did not use this method. Our models were only fine-tuned on the datasets included in the SuperGLUE benchmark.

### v4

@@ 4.3 Proximal Policy Optimization for LLM Fine-tuning on NLU Tasks
We utilize PPO to fine-tune the LLM on NLU tasks, following the training protocol outlined in section 3. The reward function is specifically designed for each NLU task. In this work, we use a simple reward function, where a reward is assigned at the end of the generation based on alignment with the ground truth labels. We use regular expression matching to extract answers from the LLMs outputs by first locating the text within the ‘<Judgement></Judgement>’ tags. Depending on the task, we then search for task-specific keywords (such as “yes”, “no”, “acceptable”, or “not acceptable”) to identify the answer. These extracted answers are compared with the ground truth to determine the appropriate rewards.
@@ 4.4 Low-Rank Adaptation
To mitigate the computational cost of full-model fine-tuning, we employ LoRA (Hu et al., 2021) during both the supervised fine-tuning and PPO stages. Instead of updating the entire model, we restrict the updates to LoRA layers, which significantly reduces the number of trainable parameters by decomposing the weight matrices into low-rank matrices.
@@ 5.1 Experimental Setup
We trained and evaluated our models on the GLUE(Wang et al., 2019) and SuperGLUE(Wang et al., 2020) benchmarks. All experiments were conducted using instruction-tuned LLAMA2-7B models(Touvron et al., 2023a) [^1] https://huggingface.co/daryl149/llama-2-7b-chat-hf. We perform both single task and multi-task fine-tuning: 1) Single-task Fine-tuning: For each subtask within GLUE and SuperGLUE, a separate task-specific LoRA module was trained independently. 2) Multi-task Fine-tuning: In the multi-task setting, datasets from different subtasks within each benchmark were combined, and a single LoRA module was trained to handle all tasks simultaneously. Please refer to Appendix A for detailed hyperparameter settings.
@@ 5.2 Baselines
  • Encoder-only models: We compare our results with encoder-only transformer models, specifically BERT-base (110M parameters) and BERT-large (340M parameters)(Devlin et al., 2019).
@@
  • Few-shot prompting: In this setting, the model is provided with both the task-specific prompt and one to five labeled examples (which ever gave the best performance) from the training dataset as demonstrations. These examples are provided as reference to guide the model in generating more accurate responses for the input query. Similarly, no task-specific fine-tuning is performed.
@@ 5.3 Results on GLUE Benchmark
Table 1: GLUE test results are scored by the evaluation server (GLUE benchmark). Average column indicates the averaged performance across all the datasets excluding the WNLI and AX datasets. F1 scores are reported for QQP and MRPC, Spearman correlations for STS-B, Matthew’s correlations for CoLA, and accuracy scores for the other tasks. Zero-shot prompting refers to prompting with task-specific prompts and an input query, while Few-shot prompting refers to prompting with task-specific prompts, 1-5 demonstrations (chosen based on the best performance), and an input query. PPO stands for proximal policy optimization, and SFT refers to Supervised Fine-tuning. “ST” represents Single-task, while “MT” represents Multi-task. The bolded results indicate the best results, and the underlined results indicate the second-best results.
In this section, we present our experiments on the GLUE benchmark, comparing the results with encoder-only models such as BERT(Devlin et al., 2019). We use the LLAMA2-7B-chat-hf model as the LLM for our evaluations. The baselines include zero-shot prompting and few-shot prompting. For fine-tuning methods, we compare both supervised fine-tuning and PPO across single-task and multi-task settings. The results are summarized in Table 1. From the results, we make the following observations.
@@
First, we observed that zero-shot prompting of the LLAMA2-7B-chat-hf model with task-specific prompts consistently underperformed compared to the smaller BERT-base model. LLAMA2-7B-chat-hf struggled notably on simpler tasks like SST-2, which only required classifying sentiment as positive or negative. This underscores the model’s weak language understanding capabilities, with zero-shot prompting proving inadequate compared to BERT-base. Second, few-shot prompting showed improvements over the zero-shot baseline, achieving an average score of 58.7 compared to 46.1, but it still lagged significantly behind the BERT-base model’s score of 79.6. Third, supervised fine-tuning (SFT) using LoRA modules for each task further boosted performance, bringing it closer to BERT’s level with an average score of 78.5, though still slightly behind BERT-base’s 79.6. Fourth, fine-tuning with PPO delivered the best results, achieving an average score of 84.6, surpassing even BERT-large’s 82.1. Moreover, zero-shot and few-shot prompting of LLAMA2-7B-chat-hf displayed a noticeable output imbalance, with a tendency to favor certain classes or values. In contrast, models fine-tuned with PPO showed no significant bias. Fifth, the total computational time for PPO is approximately 1.32 times that of SFT, indicating only a marginal increase in computational costs.
@@
Additionally, we compared the results with multi-task training, where a single LoRA module was trained across all datasets using both SFT and PPO to reduce time complexity. We found that SFT on individual tasks outperformed its multi-task fine-tuning counterpart. However, while PPO on multi-task training did not perform as well as PPO on single-task training, it still outperformed BERT-large in average performance, achieving a score of 82.9 compared to BERT-large’s 82.1. These results demonstrate that while single-task fine-tuning yields the best performance, multi-task training with PPO can still achieve competitive results, even surpassing state-of-the-art models like BERT-large.
5.4 Evaluating Zero-Shot Generalization of PPO Fine-Tuned Models and Comparison with GPT-4o
Tasks LLAMA2-7B PPO-ST LLAMA2-13B PPO-ST GPT-4o
Sentiment Analysis
Financial PhraseBank 97.2 97.7 97.5
Labelled Financial News 70.2 72.3 67.8
Mental Health 67.2 66.6 59.9
Emotion 78.0 76.4 77.6
Natural Language Inference
Babi-nli 68.3 69.4 63.2
SIGA-nli 46.2 46.3 35.4
Average 71.2 (4.3$\uparrow$) 71.5 (4.6$\uparrow$) 66.9
Table 2: Accuracy of different models across downstream tasks. For sentiment analysis tasks, models are fine-tuned on SST-2 and zero-shot evaluated on Financial PhraseBank (Malo et al., 2014), Labelled Financial News (Sood, 2024), Mental Health (Gaes, 2023), and Emotion (Saravia et al., 2018). Similarly, for natural language inference tasks, models are fine-tuned on MNLI and zero-shot evaluated on Babi-nli (Weston et al., 2015) and SIGA-nli (Nizamani et al., 2024). PPO-ST represents fine-tuning using Proximal Policy Optimization. Gains over GPT-4o model in the average row is indicated with green arrows.
We evaluate the zero-shot generalization capabilities of LLAMA2 7B and 13B models fine-tuned using PPO on a single dataset and subsequently tested across multiple other datasets (Table 2). For sentiment analysis tasks, the models were fine-tuned on SST-2 and evaluated on diverse datasets, including Financial PhraseBank (Malo et al., 2014), Labelled Financial News (Sood, 2024), Mental Health (Gaes, 2023), and Emotion (Saravia et al., 2018). Similarly, for natural language inference (NLI) tasks, the models were fine-tuned on MNLI and evaluated on Babi-nli (Weston et al., 2015) and SIGA-nli (Nizamani et al., 2024).
Our results demonstrate that PPO fine-tuning improves the zero-shot performance of LLAMA2-chat-hf models compared to GPT-4o, a strong baseline. For sentiment analysis, LLAMA2-13B-chat-hf achieves 97.7% accuracy on Financial PhraseBank, slightly outperforming GPT-4o (97.5%). On Labelled Financial News, LLAMA2-13B-chat-hf records 72.3%, exceeding GPT-4o by 4.5%. Similarly, on the Mental Health dataset, LLAMA2-7B-chat-hf achieves 67.2%, marking a notable gain of 7.3% over GPT-4o. For the Emotion dataset, LLAMA2-7B-chat-hf achieves 78.0%, with a smaller gain of 0.4%. For NLI tasks, LLAMA2-13B-chat-hf achieves 69.4% accuracy on Babi-nli, surpassing GPT-4o by 6.2%. Additionally, LLAMA2-13B-chat-hf achieves 46.3% accuracy on SIGA-nli, outperforming GPT-4o by more than 10%. On average, both 7B and 13B versions of PPO fine-tuned LLAMA2-chat-hf models demonstrate a performance gain of over 4% compared to GPT-4o, which is significantly larger in size and highly optimized.
To ensure robust comparisons, we quantify uncertainty in our evaluations by generating 100 predictions for each example in the dataset. The evaluation metric is then computed over the entire dataset for each set, yielding a distribution of values. The 95% confidence interval is defined by the 2.5th and 97.5th percentiles of this distribution. Results are presented in Table 5.
These results demonstrate the effectiveness of simple PPO fine-tuning on a single task-specific dataset in significantly enhancing model performance on similar tasks. LLAMA2-chat-hf models fine-tuned with PPO consistently outperform GPT-4o across diverse downstream tasks, reinforcing PPO fine-tuning as a robust approach for improving the NLU capabilities of LLMs.
We measured inference time on the Financial PhraseBank dataset with a batch size of 4. The BERT-base model, with 110M parameters, required 0.035s per step, while the LLAMA2-7B model, with 7B parameters and multi-token generation, took 0.997s per step. This difference is expected given the larger model size and the need for multiple forward passes in LLAMA2-7B. While LLM inference is slower, our focus is on improving natural language understanding with PPO, which achieves strong performance gains on both in-distribution and out-of-distribution NLU and NLI tasks.
5.5 Evaluation of Instruction-Following in Out-of-Distribution Tasks
To assess the instruction-following capabilities of LLMs in tasks differing from their fine-tuned format, we conduct evaluations using the LLAMA2-7B-chat-hf model fine-tuned on the SST-2 dataset. Specifically, we evaluate the performance of this model on the Amazon review task, which requires generating an integer rating between 1 and 5 based on the provided textual review. Although SST-2 and Amazon reviews both involve sentiment analysis, the two tasks differ distinctly in their input-output formatting, providing a clear measure of instruction-following adaptability.
We compare three versions of the LLAMA2-7B-chat-hf model: the original non-fine-tuned model, a version fine-tuned using SFT, and another fine-tuned with PPO. The 95% confidence intervals (CI) reported here are defined by the 2.5th and 97.5th percentiles of the bootstrap distribution. Using a consistent prompt across models, we find that the PPO-fine-tuned model achieves an accuracy of 39.35% (95% CI: 38.39, 40.29), significantly outperforming the original model, which achieves 27% accuracy (95% CI: 19.00, 36.03). Conversely, the SFT-fine-tuned model demonstrates extremely poor performance, achieving less than 1% accuracy.
Method Accuracy 95% CI
Original 27.00 (19.00, 36.03)
SFT 0.00961 (0.00, 0.03)
PPO 39.35 (38.39, 40.29)
Table 3: Performance of LLAMA2-7B-chat-hf on the Amazon Review dataset. Best results are highlighted in bold.
Qualitative analysis of sampled outputs reveals that the PPO-fine-tuned model reliably adheres to the instruction format and generates detailed reasoning to support its predictions. In contrast, the SFT-fine-tuned model often fails to adapt its responses to the required format, demonstrating limited generalization capabilities. PPO fine-tuning maintains proximity to the original model distribution via a clipping mechanism, thus preserving and enhancing the model’s intrinsic instruction-following capabilities. In contrast, SFT fine-tuning appears to narrow the model’s learned distribution to task-specific training data, negatively impacting its original instruction-following proficiency. We also evaluate the impact of fine-tuning on language modeling ability in Appendix E.
5.6 Performance Comparison Across Different LLMs
Models STS-B COPA
BERT-large 86.5 70.6
LLAMA2-7B-chat-hf
Zero-shot prompting 27.5 57.0
Few-shot prompting 45.5 73.4
PPO-ST 92.6 88.6
Qwen2.5-7B-Instruct
Zero-shot prompting 83.7 96.6
Few-shot prompting 87.0 96.0
PPO-ST 92.2 97.0
MPT-7B-chat
Zero-shot prompting 19.7 57.4
Few-shot prompting 21.7 57.2
PPO-ST 89.3 84.0
Table 4: Performance comparison of LLAMA2-7B-chat-hf, Qwen2.5-7B-Instruct(Hui et al., 2024), and MPT-7B-chat(MosaicML, 2023) models on the GLUE STS-B and SuperGLUE COPA tasks under zero-shot prompting, few-shot prompting, and PPO based fine-tuning. Results are sourced from the official GLUE benchmark and SuperGLUE benchmark evaluation servers. For STS-B, we report Spearman correlation, and for COPA, accuracy is used as the evaluation metric.
To assess the consistency of our findings across different models, we evaluated Qwen2.5-7B-Instruct and MPT-7B-chat alongside LLAMA2-7B-chat-hf on the STS-B dataset from the GLUE benchmark and the COPA dataset from the SuperGLUE benchmark. The results confirm that PPO-based fine-tuning consistently outperforms the BERT-large model, as well as the zero-shot and few-shot prompting baselines for all LLMs, highlighting its effectiveness across different LLMs. Additionally, the effect of few-shot prompting on COPA performance varies across different LLMs, indicating that different LLMs have varying capabilities to process and follow long-context instructions, which results in variable performance outcomes.
6 Conclusion
Prompting-based approaches, including zero-shot and few-shot prompting, are commonly used to adapt LLMs to downstream tasks. However, our experiments show that when applied to LLAMA2-7B-chat-hf, these methods underperform on NLU benchmarks such as GLUE and SuperGLUE (table 6), often trailing smaller encoder-only models like BERT-base. To address this, we investigate two fine-tuning strategies that update only LoRA layers for computational efficiency: SFT and PPO. While SFT yields modest improvements, PPO provides substantial gains by framing NLU tasks as reinforcement learning problems. PPO-tuned models not only outperform strong baselines like BERT-large but also generalize well across model families and tasks. Notably, PPO-trained LLAMA2-7B-chat-hf outperforms GPT-4o by 10.8% on SIGA-nli and 7.3% on the Mental Health dataset, demonstrating strong zero-shot generalization from single-task fine-tuning. More broadly, we highlight a promising direction: adapting LLMs to new tasks without labeled data by using reward-driven learning. With a well-defined reward function, PPO can steer models toward high-reward behaviors—offering a scalable, label-efficient alternative to SFT.
Limitations
This work takes an initial step toward framing NLU as a reinforcement learning problem for LLMs under 14B parameters. While our long-term goal is to reduce reliance on curated datasets by leveraging richer, task-specific reward models, we currently adopt a simple binary reward signal based on exact label matching. This design enables a controlled evaluation of PPO as an effective adaptation strategy, showing consistent gains over prompting and supervised fine-tuning. Although we present a preliminary exploration of model-driven reward functions in Appendix G, further research is needed to develop robust and generalizable reward signals that can support more complex or weakly supervised tasks without requiring extensive manual annotation. Overall, our findings suggest that casting nuanced tasks as reinforcement learning problems—through the design of appropriate environments and reward functions—offers a scalable and flexible alternative to standard fine-tuning, particularly when the model is already well-initialized through pretraining.
Appendix A Hyperparameter Settings
For PPO-based fine-tuning, grid search is performed to select the batch size in 4, 8, 12, and 16 for each task. A batch size of 24 was used across all tasks during supervised fine-tuning (SFT). The PPO epoch is set to 4, meaning that each sampled batch is used for updating the model four times. The initial learning rate for all tasks was set to $9\times 10^{-6}$. We utilized the Adafactor optimizer for PPO training and AdamW for SFT. A cosine annealing learning rate scheduler with a warmup phase was employed, where the learning rate was gradually increased during the first 10% of training steps and then reduced to one-tenth of the initial value by the end of training. We use a rank $r=16$ for the LoRA layers. We trained both PPO and SFT models until convergence on the validation set. The best hyperparameters were selected based on performance on the validation set. The final reported results for the GLUE and SuperGLUE are from their corresponding evaluation server. For evaluation, multinomial sampling with a temperature of 1 was used to generate a single response per data sample. The model generated responses with lengths between 12 and 32 tokens, with the generation process concluding using a special identifier ‘‘</Judgement>’’.
Tasks LLAMA2-7B PPO-ST LLAMA2-13B PPO-ST GPT-4o
Sentiment Analysis
Financial PhraseBank (96.2, 98.1) (96.9, 98.5) (96.6, 98.4)
Labelled Financial News (66.1, 74.6) (69.0, 76.6) (63.2, 72.2)
Mental Health (66.6, 67.7) (66.0, 67.1) (59.3, 60.5)
Emotion (77.4, 78.6) (75.8, 77.0) (77.0, 78.2)
Natural Language Inference
Babi-nli (64.3, 71.5) (65.1, 73.0) (58.8, 67.6)
SIGA-nli (39.0, 53.9) (40.6, 53.7) (28.5, 42.2)
Table 5: To quantify uncertainty in our evaluations, we generate 100 predictions for each example in the dataset. The evaluation metric is then computed for each set over the entire dataset, forming a distribution of values. The 95% confidence interval is defined by the 2.5th and 97.5th percentiles of this distribution. For sentiment analysis, models fine-tuned on SST-2 are evaluated in a zero-shot setting on Financial PhraseBank, Labelled Financial News, Mental Health, and Emotion datasets. For natural language inference, models fine-tuned on MNLI are zero-shot evaluated on Babi-NLI and SIGA-NLI.
Appendix B Reward Curve for PPO Fine-Tuning
Figure 2: Reward curve for multitask PPO fine-tuning of LLAMA2-7B-chat-hf on the GLUE dataset. The plot illustrates the relationship between training iterations (x-axis) and reward values (y-axis), demonstrating the effectiveness of the PPO optimization approach in improving model performance over time.
We present the reward curve from fine-tuning LLAMA2-7B-chat-hf using PPO in a multitask setting on the GLUE dataset. Figure 2 illustrates the reward values over training iterations, offering insights into the training dynamics of the model. The curve serves as a key performance metric, tracking the model’s learning progress across multiple tasks. The consistent upward trend demonstrates that PPO fine-tuning effectively improves LLAMA2-7B-chat-hf’s ability to generate task-relevant outputs.
Appendix C Results on SuperGLUE Benchmark
@@
Table 6: SuperGLUE test results are scored by the evaluation server (SuperGLUE benchmark). The experimental data for BERT-large and BERT-large++ are taken from the original SuperGLUE paper (Wang et al., 2020). The metrics used in the experiments are as follows: CB: F1 / Acc; MultiRC: F1 / Exact Match; ReCoRD: F1 / Exact Match; AXb: MCC; AXg: Gender parity score / Acc. For the remaining tasks not mentioned, accuracy (Acc) is reported. Average column corresponds to the averaged performance across all the datasets. For tasks with multiple evaluation metrics, we first compute the average of those metrics to obtain a single task score, which is then used in the overall average calculation. The bolded results indicate the best results, and the underlined results indicate the second-best results.
@@
We fine-tuned the LLAMA2-7B-chat-hf model using PPO on the SuperGLUE dataset and compared its performance against several baselines, including BERT-large, BERT-large++, and zero-shot and few-shot prompting of LLAMA2-7B-chat-hf. The term “BERT++” refers to a BERT model fine-tuned using the supplementary training on intermediate labeled-data tasks (STILTs) approach (Phang et al., 2018), where the model is first fine-tuned on related transfer tasks before being fine-tuned on SuperGLUE tasks. For example, MNLI from the GLUE benchmark(Wang et al., 2019) is used as an intermediate task for CB, RTE, and BoolQ(Wang et al., 2020). In contrast, our experiments with LLM did not use this method. Our models were only fine-tuned on the datasets included in the SuperGLUE benchmark.